OpenStudy (anonymous):

Jane is organizing a fundraiser to buy a ping-pong table for the community center. The table costs $500.00. Jane is asking contributors to pay for an equal share of the cost of the table. She already has five contributors lined up. What function models the cost per share, C, with respect to the number of new contributors, n? How many new contributors must Jane recruit in order for the cost of each share to be $20.00?

OpenStudy (anonymous):

Help me please!

OpenStudy (anonymous):

20

OpenStudy (anonymous):

500 = n * c c = 20 500 = n * 20 500 = n * 20 - 5 500/15 = n n = 33.33333 Did I do this right?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

explain how you got 20 please

OpenStudy (anonymous):

First, you divide the price, 500, and how much each person is going to donate, 20, to get 25 people that have to donate. Then, you would subtract 5 from 25 because u already have 5 donators.

OpenStudy (anonymous):

Thank you so much :D

OpenStudy (anonymous):

Your Welcome